# Towards the Pedagogical Steering of Large Language Models for Tutoring: A Case Study with Modeling Productive Failure

### v1

@@ 3.3 Algorithm
The Pedagogical Steering problem is caused by a misalignment between an LLM’s goal and tutoring goals, and by LLMs’ single- rather than multi-turn dimension. To steer the LLM towards a multi-turn pedagogical strategy, we introduce StratL, an algorithm that dynamically updates the tutoring intents after every student utterance and prompts the LLM to follow them at the single-turn level.
A visual representation of StratL is presented in Figure 1. StratL is an algorithm for the formalization given in Section 3.1. This algorithm is the succession of three procedures:
  1. I.
    Student State Tracing
  2. II.
    Tutor Intent Selection
  3. III.
    Intent Dependent Steering
@@
I. The Student State Tracing mirrors the function $\psi$. This step is a multi-label classification of the dialog history, and the latest student utterance at turn $t$. The selected labels (called features) $\mathbf{s}_{n}$ are chosen from a pre-defined set of possible features $\mathcal{S}$. We defined features that provide the most relevant information for the subsequent Intent Selection. These features broadly fall into the categories of (1) error types, (2) student requests, (3) states of conversation, and (4) emotional states. These feature categories each provide valuable information to be used in the Intent Selection. (1) helps plan error remediation (e.g. a reasoning and a computational mistake should not be dealt with equally) (Daheim et al. 2024; Wang et al. 2024), (2) is necessary as different student requests call for different reactions (e.g. if the student asks for the solution v.s. if they ask for a definition), (3) helps structure the conversation (e.g. knowing if the problem is solved) and (4) allow for proper emotional responses from the tutor (Lepper et al. 1993). The precise features we use are detailed in Appendix A.2. State tracing is also performed by an LLM prompted to identify these different features given the set of possible student states $\mathcal{S}$ and the dialog history. Implementation details such as the prompt we used are provided in Appendix A.2.
II. The tutor intent selection mirrors the function $\widetilde{\phi}$. We model tutor intent selection as a deterministic transition from the previous set of intents $\mathbf{i}_{n}$ to the next $\mathbf{i}_{n+1}$, conditioned on the output $\mathbf{s}_{n}$ of step I. This step can be represented by a transition graph that shows how to move from one intent to the other depending on the selected features. This graph defines the tutoring strategy $\widetilde{\phi}$. We collaborated with Learning Scientists to manually design a transition graph mirroring a PF tutoring strategy. The graph can be found in Appendix A.3. The transition rules can be defined by an expert, as we did, but we also encourage future work to learn them from annotated data. We used the theories of Scaffolding (Reiser 2004), Affective Cognition (Lepper et al. 1993) and Productive Failure (Sinha and Kapur 2019; Kapur and Bielaczyc 2012) to design a taxonomy of tutoring intents $\mathcal{I}$. This taxonomy is detailed in Appendix A.1. These intents are mainly separated into (1) Scaffolding intents (such as giving a hint) that help the student by restricting the solution space, and (2) Problematizing intents (such as asking to formalize an intuition) that help the student deepen their understanding by exploring the solution space.
@@
For our Productive Failure strategy, we prioritize light Scaffolding intents when the student is stuck and use Problematizing intents whenever the student shows an understanding of the problem, to encourage the free exploration of the problem and the generation of multiple RSMs. The exact Intent Selection transition graph we use is displayed in Figure 4 of Appendix A.3. A detailed explanation of the Learning Science principles we used to design our Productive Failure Intent Selection is given is Appendix A.5.
@@
III. The Intent Dependent Steering generates the prompt $p_{n+1}$ for the LLM Tutor conditioned on the selected tutoring intents $\mathbf{i}_{n+1}$. This prompt integrates each intent in $\mathbf{i}_{n+1}$ by concatenating intent-dependent prompt additions to the LLM’s base prompt. We define a one-to-one correspondence between each possible tutoring intent in $\mathcal{I}$ and a prompt addition. This correspondence is given in Appendix A.4. As depicted in Figure 1, the generated prompt $p_{n+1}$ is made of a base prompt template, followed by the grounding information, and the intent-specific prompt additions.
@@ 4 Experiments
Experimental Setup. As a case study, we use StratL to implement a prototype high school math tutor based on GPT-4o (gpt-4o-2024-08-06) and following a PF–inspired tutoring strategy. Our tutor discusses one problem per tutoring dialog and has access to its solution. The precise implementation of the three steps of StratL in our prototype, including (I) the detailed prompt of the Student State Tracing and the set of possible state features $\mathcal{S}$, (II) the transition rules for the Intent Selection and the set of possible intents $\mathcal{I}$, and (III) the correspondence between intents and prompt additions, are given in Appendix A.2. We also detail the model specifications and parameters.
@@ 4.1 Research Questions
We validate the capacity of our system to control the tutoring strategy employed by the LLM by assessing the capacity of our prototype to follow the PF paradigm. We want our system to do so without hindering the factual correctness or desirable conversational properties of LLMs such as their coherence. We formulate these goals in two research questions:
  1. RQ1. To what extent does StratL make the LLM Tutor follow Productive Failure?
@@
    1. RQ2.1. Does StratL influence the ability of the LLM to generate factually correct responses?
    2. RQ2.2. Does StratL influence the ability of the LLM to act like a human?
    3. RQ2.3. Does StratL influence the perceived usefulness of the tutor?
In our implementation, We also chose to handcraft the Intent Selection mechanism with the help of PF expert teachers, with the assumption that an LLM could not model such a complex strategy by choosing intents on its own. This is different from Wang et al. (2024) which lets an LLM choose how to react. We thus formulate a design question to test our assumption:
  1. DQ1. Can the intent selection be left to an LLM?
@@ 4.2 Dataset
We collected a dataset of 20 problems from Productive Failure research. These problems are unique because each has multiple possible suboptimal solutions (and a unique optimal one) and requires multiple interactions from students to explore their solution space. As such, they are high-quality testing problems to evaluate the pedagogical skills of a tutoring system.
We test our system on two problems from our dataset. Both are suitable for 9th-grade students. The problem Consistency is adapted from Kapur and Toh (2015) and is one of the most studied PF problems in the literature. This problem is an open-ended invention problem which asks students to create a measure of consistency (canonically measured by the variance) to determine the most consistent football player among three based on the number of goals they scored over the past years. An effective PF tutor for this problem would help the students construct as many ways to measure consistency as possible. The second problem, entitled Country, is an open-ended problem taken from Kerrigan (2018) which asks students to find a way to fairly distribute voting seats for the different states of a country based on their population. For this problem, a good PF tutor would help the student without giving away any of the crucial reasoning steps necessary to solve this problem, such as using proportionality. The exact problem statements, their corrections, and the dataset specification are provided in Appendices A.6 and A.7.
@@ 4.3 Baselines
We test our StratL-controlled tutor LLM against three ablation versions.
  • V1:
    Full StratL. An LLM controlled by our implementation of StratL.
  • V2:
    No Intent. Baseline version leaving complete freedom to the tutor LLM to choose the best way to answer without conditioning on any intent. We use it to assess the tutoring strategy fidelity of our system.
  • V3:
    Constant Intent. This version constantly uses the same intent (labeled “Seek Strategy” in Table 3 of Appendix A.1) that only asks the student to proceed with the exercise. We use this baseline to assess the importance of a well-crafted Intent Selection procedure.
  • V4:
    LLM Intent. In this version, the choice of the intent is left to an LLM similarly as in the related work (Wang et al. 2024). We use it to assess whether the intent’s choice needs to be determined by an expert-designed procedure or if a natural language description of the tutoring strategy to follow is enough (DQ1). Its prompt is given in Appendix A.10.
@@ 4.4 Evaluation Metrics
To quantitatively assess the Productive Failure fidelity of the different versions (their success in following PF), we have to rely on proxy variables. We propose as a main proxy the number of student-generated Representation and Solution Methods (RSMs, which can be thought of as partial solutions). As explained in Section 2, the diversity of student-generated RSMs is the first predictor of learning gains with PF (Kapur and Bielaczyc 2012; Kapur and Rummel 2012). We thus use this number to measure the PF fidelity of the tutor.
However, not every problem easily admits multiple RSMs. In our case, only Consistency does, as an invention problem. We thus define another proxy, the Productive Failure Score. One of the core ideas of PF is to encourage the student to construct their own understanding of the problem by exploring it by themselves, even if they do not succeed in solving it. Building on this idea, we define for each problem a grading rubric used to compute the tutor’s Productive Failure Score: For each problem, we identify a list of critical reasoning steps that represent different conceptual understandings of the problem. To assess a tutoring conversation, we give a point to the tutor for every crucial reasoning step that the student found on their own. We do not award the point if the tutor reveals it or directly hints to it. This metric can be thought of as an adjusted sub-step success rate. An efficient PF tutor would help the student explore the problem and thus achieve these conceptual understandings, without giving them any direct hint or revealing them, resulting in a high PF Score. We display the precise sub-steps for both testing problems in Appendix A.8.
@@ 4.5 Preliminary Study with Simulated Students
Because of ethical concerns (Kasneci et al. 2023), we first tested our prototype with GPT-4–simulated students before exposing it to humans. This method of using simulated students, known as interactive tutoring, is commonly employed in the literature (Sonkar et al. 2023; Macina et al. 2023a). Importantly, we do not assume that the LLM is able to accurately model students’ learning, especially given the somewhat counter-intuitive aspect of PF. However, this assumption is not necessary for our preliminary study as we are interested in the ability of our algorithm to follow a PF tutoring strategy, rather than the learning gains it produces. We simulated three conversations for each of the two problems and of the four versions of the tutor (V1–4). We use a scoring rubric to manually annotate the conversations to compute the PF Score and number of RSMs as described in Section 4.4.
@@ 4.6 Field Test with Students
We ran the field test with 17 9th-grade students from a high school in Singapore. As we explain in Section 5.1, we did not include V3 and V4 in the field test following the preliminary study results. Each student was randomly assigned to one of two conditions (V1 or V2), and they had 15 minutes to interact with the tutor. We manually annotated their conversations using our rubric to compute the PF score and the number of generated RSMs.
Post-study Questionnaire. After the tutoring session, we asked each student about their experience with the tutor. The aim was to detect any spillover effect of StratL on the other desirable properties of the tutor LLM. Specifically, students were asked to rate on a 3-point Likert scale the following three aspects of their AI tutor:
  1. 1.
    The tutor is factually coherent in its answers.
  2. 2.
    The tutor answers with empathy like a teacher.
  3. 3.
    The tutor is helpful in solving the problem.
@@ 5.1 Preliminary Study with Simulated Students
Table 1: Productive Failure fidelity metrics for different tutor versions and different problems in the preliminary study with simulated students. Best values in bold.
# RSMs PF.S
@@
DQ1.
Using an LLM prompted with a natural language description of the tutoring strategy is not enough for the Intent Selection. The preliminary study in Table 1 suggests that StratL (V1) is successful in steering the LLM tutor towards PF, as evidenced by the greater number of RSMs elicited by this version. We observe that the version using StratL with an LLM responsible for the Intent Selection (V4) did perform better than the baseline LLM (V2), but not as well as StratL with our handcrafted Intent Selection mechanism. This result contributes to an answer to DQ1 by showing that providing a natural language description of the tutoring strategy to follow and letting an LLM choose the intents accordingly does not yield satisfactory results in terms of tutoring strategy fidelity.
RQ1.
The strategy followed by an LLM controlled by StratL is determined by the Intent Selection procedure. Overall, V3 achieves a rather high PF Score on the two problems, since it never provides any hint to students. However, this result is more a consequence of the way we defined the PF score rather than a sign of PF fidelity, as V3 does not perform as well on our second proxy: It does not manage to elicit as many RSMs from students as V1, with an average of only 1.34. This shows that the strategy followed by an LLM controlled by StratL is indeed determined by the Intent Selection procedure as it is the only part that differs between V1 and V3.
Based on these results, we decided not to include V3 and V4 in our field test with real students to focus on the best-performing version (V1) while keeping the non-StratL baseline (V2) for comparison.
Accuracy of Student State Tracing
The functioning of our system as a whole is contingent upon the functioning of the Student State Tracing. To test its reliability, we generate ten tutoring conversations for both problems and one author manually annotates them with ground truth state features $\mathbf{s}_{t}$. We compare the outputs of the Student State Tracing with the annotated data. As a multi-label classification task, we compute for each possible label (state feature) the precision, recall, and micro F1 score. We compare three different implementations of the State Tracing step and provide more details about our methodology and results in Table 6 of Appendix A.9. Our simple procedure already achieves satisfactory results of 0.77 for the micro F1 score. The Student State Tracing is a task of its own, and we encourage researchers to try different classification models to improve the reliability of this step.
Figure 2: Results from the field test and their confidence intervals. ‘PF.S’ refers to the ‘Productive Failure Score’ as defined in Section 4.4 and is out of 4. ‘RSMs’ denotes the number of student-generated RSMs. StratL (V1) succeeds in steering the LLM to follow Productive Failure as compared to V2.
@@ 5.2 Field Test with Students: Measuring end-to-end Effectiveness
RQ1. Tutoring Strategy Fidelity
StratL is successful in steering the LLM to follow the Productive Failure–inspired tutoring strategy. We report in Figure 2 the results of the protocol defined in Section 4.4. We note that the PF score (out of 4) is used as a proxy for the extent to which the tutor followed the PF tutoring strategy. We also report the average number of student-generated RSMs (partial solutions) on the problem Consistency. This metric is our best proxy for PF fidelity, as one of the goals of the PF tutoring strategy is to maximize this number, as explained in Section 2. Out of the two test problems, the number of RSMs can only be measured on Consistency as it is the only invention problem.
Table 2: Feedback questionnaire results (out of 3). ‘N’ refers to the number of students who answered the questionnaire, ‘Factual’ to the factual correctness of the tutor, ‘Human’ to its ability to communicate in natural language like a human, and ‘Helpful’ to its helpfulness as perceived by the student. We do not detect any harmful effect of StratL (V1) on these properties.
Consistency Country
N Factual Human Helpful N Factual Human Helpful
V1. StratL 5 1.80 1.60 1.20 3 3.00 2.67 2.67
V2. No Intent 4 2.25 1.75 2.00 5 3.00 2.20 2.80
@@
RQ2. Student Feedback Questionnaire
We report in Table 2 the mean feedback questionnaire results for each problem on a 3-point Likert scale).
@@
RQ2.1 StratL has no or little influence on the capacity of the tutor LLM to generate factually coherent responses. StratL (V1) scores similarly or slightly less than an LLM with no intent (V2) on factual correctness across the two problems. We do not detect any statistically significant difference for any of the versions.
@@
RQ2.2 StratL does not influence the capacity of the LLM to generate human-like answers. None of the two versions is perceived as significantly more ‘human-like’ than the other. StratL does not seem to impede the ability of an LLM to produce natural language.
@@
RQ2.3 StratL can influence the perceived helpfulness of the tutor LLM, through the tutoring strategy it makes it follow. We find that the baseline with no intents (V2) is perceived by the students as more helpful than V1. This is not surprising, as the PF tutoring strategy can be frustrating or misunderstood by students. Indeed, the PF strategy is designed to only give help when it is strictly necessary, which means that students spend more time being stuck and exploring suboptimal solutions. This can lead students to perceive the tutor as less helpful compared to a version that would give more help.
5.3 Qualitative Insights from Students Feedback

### v2

@@ 3.2 Algorithm
To steer the LLM towards a multi-turn pedagogical strategy, we introduce StratL, an algorithm for the formalization (3) given in Section 3.1. StratL dynamically updates the tutoring intents after every student utterance and prompts the LLM to follow them at the single-turn level. A visual representation of StratL is presented in Figure 1. This algorithm is based on three sub-procedures: I. Student State Tracing, II. Tutor Intent Selection and III. Intent Dependent Steering.
@@
I. The Student State Tracing mirrors the function $\psi$. This step is a multi-label classification of the dialog history and latest student utterance at turn $t$. The selected features $\mathbf{s}_{n}$ are chosen from a pre-defined set of possible features $\mathcal{S}$. These features fall into four categories: (1) mistake types, (2) student requests, (3) conversation states, and (4) emotional states. These categories each provide valuable information to be used in the Intent Selection, i.e.: (1) helps plan error remediation (e.g. a reasoning and a computational mistake should not be dealt with equally) Daheim et al. (2024); Wang et al. (2024a), (2) is necessary as different student requests call for different reactions (e.g. if the student asks for the solution v.s. if they ask for a definition), (3) helps structure the conversation (e.g. knowing if the problem is solved) and (4) allow for proper emotional responses from the tutor Lepper et al. (1993). The precise features we use are detailed in Appendix B.2.1. State tracing is performed by an LLM prompted to identify these different features given the set of possible student states $\mathcal{S}$ and the dialog history. Implementation details such as the prompt we used are provided in Appendix B.2.1.
II. The Tutor Intent Selection mirrors the function $\widetilde{\phi}$. We model tutor intent selection as a deterministic transition from the previous set of intents $\mathbf{i}_{n}$ to the next set of intents $\mathbf{i}_{n+1}$, conditioned on the output $\mathbf{s}_{n}$ of step I. The transition rules are defined by a learning scientist expert. We represent these transitions with a graph. Please note that, unlike traditional deterministic state transition models, our graph allows to transition from a set of nodes to another by following multiple concurrent transitions.
@@
We used the theories of Scaffolding Reiser (2004), Affective Cognition Lepper et al. (1993), and PF Sinha and Kapur (2019); Kapur and Bielaczyc (2012) to design a taxonomy of tutoring intents $\mathcal{I}$. These intents are mainly separated into: (1) scaffolding intents, helping the student by restricting the solution space (e.g. giving a hint), and (2) problematizing intents, helping the student deepen their understanding by exploring the solution space (e.g. asking to formalize an intuition). For our PF strategy, we prioritize scaffolding when the student is stuck and use problematizing intents whenever the student shows an understanding of the problem, to encourage the free exploration of the problem and the generation of multiple RSMs. The exhaustive taxonomy is displayed in Table 5 of Appendix B.1 and a detailed explanation of the Learning Science principles we used to design our PF Intent Selection is given in Appendix B.3.
Our PF Intent Selection graph is represented in Figure 2. Each node represents an intent of $\mathcal{I}$. Arrows correspond to transitions between intents. We initialize start at the ‘Ask for the Next Step’ node and move at each tutoring turn. The letters used in the logical conditions on the arrows are codes corresponding to Student State Tracing features (detailed in Figure 8 of Appendix B.2.1). For the sake of clarity, transitions that can be taken from any intent are represented by arrows with no starting node.
@@
III. The Intent Dependent Steering generates the LLM Tutor prompt $p_{n+1}$ conditioned on the selected tutoring intents $\mathbf{i}_{n+1}$. We defined a one-to-one correspondence between the intents in $\mathcal{I}$ and short prompt chunks. The precise correspondence is given in Appendix B.2.2.
@@ 4 Experiments
Experimental Setup. As a case study, we use StratL to implement a prototype high school math tutor based on GPT-4o (gpt-4o-2024-08-06) and following a PF–inspired tutoring strategy. Our tutor discusses one problem per tutoring dialog and has access to its solution. The exact prompts we used in our prototype, and the exhaustive list of state features, intent taxonomy, and intent-to-prompt correspondences are given in Appendix B.2.
@@ 4.1 Research Questions
We validate the ability of our system to control the tutoring strategy employed by the LLM by assessing the capacity of our prototype to follow the PF paradigm. Ideally, without hindering the factual correctness or desirable conversational properties of LLMs such as their coherence. We formulate these goals in two research questions:
  1. RQ1. To what extent does StratL make the LLM Tutor follow PF?
@@
We chose to handcraft the Intent Selection mechanism with the help of PF experts. This is different from Wang et al. (2024a) which lets an LLM choose how to react. We thus formulate a design question to test our assumption: DQ1. How does the intent selection by LLM compare to our approach?
@@ 4.2 Dataset
We compiled a dataset of 15 PF problems. These problems are open-ended, ill-structured, and engaging. They feature multiple suboptimal solutions (and a single optimal one) and require multiple interactions from students to explore their solution space. These serve as high-quality testing problems for evaluating a tutoring system’s pedagogical skills. Our algorithm is tested on two 9th-grade math problems (successfully solved by 62% of our students in the field test). The first problem, Consistency, is adapted from Kapur and Toh (2015) and is a well-studied invention problem where students devise a consistency measure (typically variance) to compare football players’ performance. An effective PF tutor for this problem would help the students construct as many ways to measure consistency as possible. The second problem, entitled Country, is an open-ended problem taken from Kerrigan (2018) requiring students to fairly allocate voting seats across states based on their population. On this problem, an effective PF tutor would guide reasoning without explicitly providing key steps like the idea of using proportionality. The exact problem statements, solutions, and the dataset specifications, intended use, and license are provided in Appendices A.1 and A.2.
@@ 4.3 Baselines
We test our StratL-controlled tutor LLM (V1) against three ablation versions. V2: No Intent is a baseline LLM prompted to act like a tutor with our base prompt (but no intent-dependent prompt). V3: Constant Intent always uses the same intent (“Ask for Next Step” in Table 5 of Appendix B.1) that only asks the student to proceed with the exercise. We use this baseline to assess the importance of the Intent Selection procedure. V4: LLM Intent use an LLM to perform the Intent Selection (with the prompt given in Appendix B.2.3Wang et al. (2024a), and we use it to answer DQ1.
@@ 4.4 Evaluation Metrics
To quantitatively assess the PF fidelity of the different versions (their success in following PF), we have to rely on proxy variables. We propose as a main proxy the number of student-generated Representation and Solution Methods (RSMs), which can be seen as the different ways students represent the problem and try to solve it. As explained in Section 2.3, the diversity of student-generated RSMs is the first predictor of learning gains with PF Kapur and Bielaczyc (2012); Kapur and Rummel (2012). We thus use this number to measure the PF fidelity of the tutor.
However, not every problem easily admits multiple RSMs. In our case, only Consistency does. We thus define another proxy, the PF Score. One of the core ideas of PF is to encourage the student to construct their own understanding of the problem by exploring it by themselves, even if they do not succeed in solving it. Building on this idea, we define for each problem a grading rubric used to compute the tutor’s PF Score: For each problem, we identify a list of critical reasoning steps that represent different conceptual understandings of the problem. To assess a tutoring conversation, we give a point to the tutor for every crucial reasoning step that the student found on their own. We do not award the point if the tutor reveals it or directly hints to it. This metric can be thought of as an adjusted sub-step success rate. An efficient PF tutor would help the student explore the problem and thus achieve these conceptual understandings, without giving them any direct hint or revealing them, resulting in a high PF Score. We display the precise sub-steps for both testing problems in Appendix A.3.
@@ 4.5 Preliminary Study with Simulated Students
Due to ethical concerns Kasneci et al. (2023), we first tested our prototype with GPT-4–simulated students before involving humans. This method of using simulated students, known as interactive tutoring, is used in the literature Sonkar et al. (2023); Macina et al. (2023a). Importantly, we do not assume that the LLM can accurately model students’ learning. However, this assumption is not necessary for our preliminary study as we are interested in the ability of our algorithm to follow a PF tutoring strategy, rather than the learning gains it produces. We simulated three conversations for each of two problems across four tutor versions (V1–4) and used a scoring rubric to annotate conversations, computing the PF Score and RSM count (as described in Section 4.4).
@@ 4.6 Field Test with Students
We ran the field test with 17 9th-grade students from a high school in Singapore. Each student was randomly (uniformly) assigned to one of two conditions (V1 or V2, chosen as described in Section 5.1) and one of the two exercises. They had 15 minutes to interact with the tutor. We manually annotated their conversations using our rubric to compute the PF score and the number of generated RSMs.
Post-study Questionnaire. After the tutoring session, students were asked about their experience to assess potential spillover effects of StratL on the tutor LLM’s desirable properties. Using a 3-point Likert scale, they rated the AI tutor on: a) coherence, b) empathy, and c) helpfulness.
@@ 5.1 Preliminary Study with Simulated Students
Table 1: PF fidelity metrics for different tutor versions and different problems in the preliminary study with simulated students. Best values in bold.
# RSMs PF Score
@@
DQ1. Using an LLM prompted with a natural language description of the tutoring strategy is not enough for the Intent Selection. The preliminary study in Table 1 suggests that StratL (V1) is successful in steering the LLM tutor towards PF, as evidenced by the greater number of RSMs elicited by this version. We observe that the version using StratL with an LLM responsible for the Intent Selection (V4) did perform better than the baseline LLM (V2), but not as well as StratL with our handcrafted Intent Selection mechanism. This result contributes to an answer to DQ1 by showing that providing a natural language description of the tutoring strategy to follow and letting an LLM choose the intents accordingly does not yield satisfactory results in terms of tutoring strategy fidelity.
RQ1. The strategy followed by an LLM controlled by StratL is determined by the Intent Selection procedure. Overall, V3 achieves a rather high PF Score on the two problems, since it never provides any hint to students. However, this result is more a consequence of the way we defined the PF score rather than a sign of PF fidelity, as V3 does not perform as well on our second proxy: It does not manage to elicit as many RSMs from students as V1, with an average of only 1.34. This shows that the strategy followed by an LLM controlled by StratL is determined by the Intent Selection procedure as it is the only part that differs (V1 vs. V3).
Based on our preliminary answers to DQ1 and RQ1, we decided not to include V3 and V4 in our field test with real students to focus on the best-performing version (V1) while keeping the non-StratL baseline (V2) for comparison.
Accuracy of Student State Tracing The functioning of StratL as a whole is contingent upon the functioning of the Student State Tracing. To test its reliability, we generate ten tutoring conversations for both problems and the first author manually annotates them with ground truth state features $\mathbf{s}_{t}$. We compare the outputs of the Student State Tracing with the annotated data. As a multi-label classification task, we compute for each state feature the precision, recall, and F1 score. We compare three different implementations of the State Tracing step and provide more details about our methodology and results in Table 4 of Appendix A.4. Our implementation achieves a satisfactory micro F1 score of $0.77$. The Student State Tracing is a task of its own, and we encourage researchers to try different classification models to improve its reliability.
@@ 5.2 Field Test with Students: Measuring end-to-end Effectiveness
Figure 3: Field test results and their confidence intervals. ‘PF.S’ refers to the ‘PF Score’ as defined in Section 4.4 and is out of 4. ‘RSMs’ denotes the number of student-generated RSMs. StratL (V1) succeeds in steering the LLM to follow PF as compared to V2.
Table 2: Feedback questionnaire results (out of 3). ‘N’ refers to the number of answers to the questionnaire. We do not detect any harmful effect of StratL (V1) on the LLM tutor’s coherence, empathy, or perceived helpfulness.
N Coherence Empathy Helpful.
Problem: Consistency
V1. 5 1.80 1.60 1.20
V2 4 2.25 1.75 2.00
Problem: Country
V1 3 3.00 2.67 2.67
V2. 5 3.00 2.20 2.80
RQ1. Tutoring Strategy Fidelity StratL is successful in steering the LLM to follow the PF tutoring strategy. We report in Figure 3 the results of the protocol defined in Section 4.4. We note the PF score (out of 4) is used as a proxy for the extent to which the tutor followed the PF tutoring strategy. We also report the average number of student-generated RSMs (partial solutions) on the problem Consistency. This metric is best proxy for PF fidelity, as one of the goals of the PF tutoring strategy is to maximize this number, as explained in Section 2.3. Out of the two test problems, the number of RSMs can only be measured on Consistency as it is the only invention problem.
@@
RQ2. Student Feedback Questionnaire The feedback questionnaire did not reveal particular spillover effects of StratL on the coherence, empathy or perceived helpfulness of the tutor LLM.
@@
StratL has no or little influence on the capacity of the tutor LLM to generate coherent responses. StratL (V1) scores similarly or slightly less than an LLM with no intent (V2) on coherence across the two problems: $1.80$ out of $3$ for (V1) vs. $2.25$ for (V2) when used on problem Consistency, and $3.00$ for both on Country. We do not detect any statistically significant difference.
@@
StratL does not influence the capacity of the LLM to generate empathetic answers. None of the two versions is perceived as significantly more empathetic than the other ($1.60$ for V1 vs. $1.75$ for V2 on Consistency and $2.67$ vs. $2.20$ on Country).
@@
StratL can influence the perceived helpfulness of the tutor LLM, through the tutoring strategy it makes it follow. We find that the baseline with no intents (V2) is perceived by the students as more helpful than V1 ($2.00$ versus $1.20$ respectively on Consistency and $2.80$ versus $2.67$ on Country). This is not surprising, as the PF tutoring strategy can be frustrating or misunderstood by students. Indeed, the PF strategy is designed to only give help when it is strictly necessary, which means that students spend more time being stuck and exploring suboptimal solutions. This can lead students to perceive the tutor as less helpful compared to a version that would give more help. We report the detailed results in Table 2.
5.3 Questionnaire Feedback